# Landscape of Correlated Orders in Strained Bilayer Nickelate Thin Films

### v2

@@ Abstract
The discovery of high-temperature superconductivity in bilayer nickelates La3Ni2O7 under pressure has sparked significant research interest. This interest has been further fueled by the recent achievement of superconductivity in compressed thin films at ambient pressure, although the origin and underlying mechanism remain elusive. In this work, we explore the electronic structures and instabilities of strained thin films on substrates to identify the key factors for achieving superconductivity, using first-principles and functional renormalization group (FRG) calculations. Our findings suggest that the compressed NiO2 bilayer at the interface is unlikely to exhibit superconductivity due to the elongation of outer apical Ni-O bond length and electron-doping effect. In contrast, the NiO2 bilayer slightly away from the interface shows density-wave instability when undoped or slightly hole-doped. However, when this bilayer is moderately hole-doped, leading to the emergence of a hole pocket around the M point, density-wave instability is suppressed and the system exhibits robust $s_{\pm}$-wave superconductivity, which may account for superconductivity observed in thin films. For the stretched NiO2 bilayer, robust spin-density-wave instability is observed due to enhanced Fermi surface nesting, despite the presence of a hole pocket around the M point. Potential experimental implications are discussed. Our study highlights the crucial role of fermiology in determining electronic instability and establishes a unified scenario for superconductivity in both pressurized bulk and strained thin films of bilayer nickelates.
@@
Despite decades of extensive research efforts into nickelates, the nickel age of superconductors only began in 2019 with the discovery of superconductivity in ”infinite-layer” nickelates, specifically (Sr,Nd)NiO2 thin films grown on a substrate Li et al. (2019); Osada et al. (2020); Pan et al. (2022); Wang et al. (2022); Ding et al. (2023). Remarkably, four years later, nickelate superconductors entered the high-temperature regime akin to cuprates with the observation of superconductivity in a novel type of bilayer nickelate La3Ni2O7 (LNO) within the Ruddlesden-Popper phase under pressure, boasting an extraordinarily high transition temperature (Tc) of approximately 80 K Sun et al. (2023). Superconductivity emerges abruptly following a structural transition from the Amam phase to the Fmmm/I4mmm phase as pressure increases. In contrast to the $d^{9}$ electronic configuration of cuprates and ”infinite-layer” nickelates, the Ni2.5+ in LNO exhibits a $d^{7.5}$ configuration, with both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals contributing to the low-energy electronic structure due to the presence of apical oxygens between layers Luo et al. (2023); Zhang et al. (2023a); Lechermann et al. (2023a); Sakakibara et al. (2024a); Gu et al. (2025); Yang et al. (2024); Liu et al. (2024). Additionally, the trilayer nickelate La4Ni3O10 exhibits superconductivity under pressure with a lower Tc of 20-30 K Sakakibara et al. (2024b); Li et al. (2024a); Zhu et al. (2024); Zhang et al. (2025); Li et al. (2024b). Both bilayer and trilayer nickelates have displayed multiple correlated phenomena, such as density wave transitions and non-Fermi liquid behaviors Sun et al. (2023); Zhang et al. (2024a); Liu et al. (2024); Chen et al. (2024); Ren et al. (2025). The pairing mechanism and symmetry are under intensive theoretical study Yang et al. (2023a); Lu et al. (2024); Oh and Zhang (2023); Gu et al. (2025); Liu et al. (2023); Qu et al. (2024); Sakakibara et al. (2024a); Yang et al. (2023b); Qin and Yang (2023); Lu et al. (2023); Tian et al. (2024); Zhang et al. (2023b, 2024b); Jiang et al. (2024a); Lechermann et al. (2023b); Liao et al. (2023); Ryee et al. (2024); Luo et al. (2024); Fan et al. (2024); Jiang et al. (2024b, c); Zhan et al. (2025); Xia et al. (2025); Bejas et al. (2025) and calculations suggest that the pressure-driven Lifshitz transition, where the $d_{z^{2}}$ interlayer bonding state crosses the Fermi level to form a hole pocket, is crucial for the emergence of bulk superconductivity Sun et al. (2023). Spin fluctuations are believed to promote an $s_{\pm}$-wave pairing following the Lifshitz transition Sakakibara et al. (2024a); Gu et al. (2025); Yang et al. (2023a); Liu et al. (2023). Other studies emphasize the importance of interlayer exchange coupling, Hund’s rule coupling, and orbital hybridization in generating interlayer $s$-wave pairing Lu et al. (2024); Oh and Zhang (2023); Qu et al. (2024). However, further experimental investigation into the electronic structures and pairing gaps in the bulk nickelates is challenged due to the high-pressure conditions required for superconductivity to occur.
@@
Intriguingly, recent studies have shown that compressively strained LNO thin films grown on the SrLaAlO4 (SLAO) substrate exhibit superconductivity with a Tc exceeding 40 K at ambient pressure Ko et al. (2025); Zhou et al. (2025); Liu et al. (2025); Bhatt et al. (2025). Although this Tc is only about half that of the bulk value, it provides a tunable and powerful platform for investigating the mechanisms underlying high-Tc superconductivity. First-principles calculations indicate that compressive strain eliminates the hole pocket around the M point, whereas tensile strain results in Fermi surfaces similar to those of the bulk material under pressure Geisler et al. (2025); Zhao and Botana (2025). However, superconductivity has not been observed in tensile-strained LNO thin films on substrates Ko et al. (2025); Bhatt et al. (2025), such as SrTiO3, which seems to contradict the proposed mechanism for superconductivity in the bulk. Recent ARPES measurements have revealed heavy hole doping in La2.85Pr0.15Ni2O7 ultrathin films, identifying a diffusive hole pocket around the M point Li et al. (2025); Yue et al. (2025). These findings motivate a comprehensive exploration of the key ingredients for superconductivity in bilayer nickelates and prompt several key questions regarding LNO thin films: (1) Which bilayer is responsible for superconductivity in LNO thin films, and what is the underlying mechanism? (2) Why do tensile-strained LNO thin films, despite hosting Fermi surfaces similar to the bulk, fail to exhibit superconductivity? (3) Do bulk and thin film LNO share the same pairing mechanism?
To address these questions, we examine the electronic structures of LNO thin films on SLAO substrate and explore competing correlated instabilities, such as superconductivity and density-wave states, for NiO2 bilayers at different locations and under different in-plane strains. We employ first-principles and functional renormalization group (FRG) calculations to conduct our analysis. Our findings reveal that the NiO2 bilayer at the interface is heavily electron-doped due to the SLAO substrate and exhibits asymmetry between two Ni sites due to the elongation of outer apical Ni-O bond at the interface. This causes a significant downshift in the onsite energy of the $d_{z^{2}}$ orbital. Our FRG analysis indicates that the instability in this NiO2 bilayer manifests as a spin density wave (SDW) over a broad range of interaction parameters due to Fermi surface nesting, suggesting that superconductivity at the interface is unlikely. For the NiO2 bilayer slightly away from the interface, an SDW is still favored when it is undoped or slightly hole-doped. However, when this bilayer is moderately hole-doped, leading to a Lifshitz transition around the M point and the appearance of a hole pocket, additional Fermi surface nesting occurs, and it results in strong competition between two SDW orders, destabilizing the SDW instability. Remarkably, both Fermi surfaces nesting occur between bonding and antibonding pockets, resulting in interlayer antiferromagnetic coupling in both SDWs. Thus, their fluctuations promote interlayer pairing, leading to a robust $s_{\pm}$-wave superconductivity. This may account for the superconductivity observed in compressed thin films. Additionally, we find that tensile strain significantly enhances Fermi surface nesting between the $\alpha$ and $\beta$ pockets as well as the $\alpha$ and $\gamma$ pockets, inducing an SDW instability despite the presence of the hole pocket around the M point. We discuss a unified scenario for superconductivity in both bulk and thin films and explore potential experimental implications.
@@
Electronic structures of LNO/SLAO interface. We start with the crystal and electronic structures of three-unit-cell (3UC) LNO on the SLAO substrate. The tetragonal LNO is assumed and initial interface is constructed according to the structure from scanning transmission electron microscope measurements Zhou et al. (2025); Yue et al. (2025) and further relaxation is done for the 3UC LNO and the first two layers of substrate. The details of first-principles calculations are shown in the supplementary materials (SM). Fig. 1(a) shows the optimized crystal structure, where only three bilayers of the film are shown along with part of the substrate, and the complete structural diagram is available in SM. The in-plane compressive strain from the substrate will generate an elongation along the c-axis for LNO, which mainly occurs on the LaO layers according to our calculations. To clearly demonstrate structural changes, we select the first and third bilayers as representative examples and show their structural parameters. The out-of-plane distance between two Ni sites $d_{\text{Ni}-\text{Ni}}$ and outer apical Ni-O bond length $d_{\text{Ni}-\text{O}}$ within these bilayers are indicated in gray and red numbers, respectively. Across these bilayers, $d_{\text{Ni}-\text{Ni}}$ is approximately 4 Å, close to the bulk value, and this value is slightly longer in the 1st bilayer than the 3rd bilayer due to the effect of substrate. Notably, the outer apical Ni-O distance is increased to 2.3 Å  away from the interface due to the elongation of LaO layers and $d_{\text{Ni}-\text{O}}$ in the 1st bilayer is elongated to 2.6 Å. This elongation of the outer apical Ni-O bond at the interface results in a significant reduction in the onsite energy of the $d_{z^{2}}$ orbital in the 1st bilayer. Besides structural modifications, there is electrostatic doping from the substrate. Fig. 1(b) displays charge density differences and blue and red indicate regions of charge accumulation and depletion, respectively, suggesting a net electron transfer from the substrate to the first bilayer.
@@
Furthermore, the orbital-resolved band structures along high-symmetry paths for the 1st and 3rd bilayers are presented in Fig. 1(c) and (d), with a focus on the $e_{g}$ orbitals ($d_{x^{2}-y^{2}}$ and $d_{z^{2}}$), as the fully occupied $t_{2g}$ orbitals lie well below the Fermi level. In the 1st bilayer, the reduced onsite energy of the $d_{z^{2}}$ orbital induces a significant downward shift of its associated bands, as seen in Fig. 1(d). This shift causes the $d_{z^{2}}$ bonding band to move further from the Fermi level, while the $d_{z^{2}}$ anti-bonding states approach and make contact with the Fermi level at the $\Gamma$ point. The electron doping of this bilayer is estimated to be about 0.34. As a result, the 1st bilayer hosts only one $\beta$ hole pocket around the M point, contributed by the mixture of $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals, and an $\alpha$ electron pocket around the $\Gamma$ point contributed by the $d_{x^{2}-y^{2}}$ orbital. The slight asymmetry between two Ni sites splits the $d_{x^{2}-y^{2}}$ bands along the $\Gamma-M$ path. The further inclusion of interaction in LDA+U formalism will induce an additional electron pocket around the $\Gamma$ point from the $d_{z^{2}}$ anti-bonding state (see SM). In contrast, the bilayer away from the interface, such as the 2nd and 3rd bilayers, shows no charge doping and its band structure closely resembles that of the bulk band structure at ambient pressure, as shown in Fig. 1(c). The $d_{z^{2}}$ bonding state just drops below the Fermi level Zhao and Botana (2025), yielding an electron pocket around the $\Gamma$ point and one hole pocket around the M point. Based on these electronic structures, we conduct model analysis and numerical calculations on correlated states to uncover the scenario of superconductivity within this heterostructure.
[…]
@@
Scenarios of superconductivity in compressed LNO thin films. Although superconductivity has been experimentally achieved in strained LNO ultrathin films, identifying which NiO2 bilayer is responsible remains elusive. Understanding this is crucial for revealing the underlying mechanism of superconductivity, as there is a significant distinction between the electronic structures of NiO2 bilayers near the interface and those further away. To investigate this issue, we construct effective tight-binding models by fitting hopping parameters to ab initio band structures and study electronic instability by varying carrier doping and interaction parameters, based on FRG calculations. As discussed earlier, the Ni $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals dominate the low-energy physics in strained LNO thin films, similar to bulk LNO at both ambient and high pressures Luo et al. (2023); Zhang et al. (2023a); Lechermann et al. (2023a); Sakakibara et al. (2024a); Gu et al. (2025); Yang et al. (2024); Liu et al. (2024). Therefore, a two-orbital tight-binding (TB) model can accurately describe the low-energy band structure Luo et al. (2023); Zhang et al. (2023a); Lechermann et al. (2023a); Sakakibara et al. (2024a); Gu et al. (2025), and the corresponding Hamiltonian is given by,
@@
Here $i,j=(m,l)$ label the in-plane lattice site ($m$) and top/bottom layer index ($l=t,b$), $\sigma$ labels spin, and $\alpha,\beta=x,z$ represent the Ni orbitals with $x$ denoting the $d_{x^{2}-y^{2}}$ and $z$ the $d_{z^{2}}$ orbital. The chemical potential $\mu$ varies with charge doping. Details of the Hamiltonian matrix are provided in the SM. Compared to the bulk case, the asymmetry between the two Ni sites in the 1st bilayer at the interface breaks the mirror symmetry connecting them, resulting in slightly different onsite energies and in-plane hopping for the two layers. The bilayer slightly away from the interface preserves this mirror symmetry and shares the same TB model with the bulk, though with modified hopping parameters and onsite energies due to strain. For the electronic interactions, we adopt the general multi-orbital Hubbard interactions,
@@
where $U/U^{\prime}$ is the intra-orbital/inter-orbital Hubbard repulsion, $J_{H}$ is the Hund ’s coupling, and $J_{P}$ is the pair-hopping interaction. The standard Kanamori relations Castellani et al. (1978) $U=U^{\prime}+2J_{H}$ and $J_{H}=J_{P}$ are adopted in this work. To resolve electronic instabilities, we employ the FRG approach, which treats all particle-hole and particle-particle channels on equal footing, providing an unbiased depiction of correlated states from weak to intermediate coupling regimes Wang et al. (2009); Metzner et al. (2012); Platt et al. (2013). As the RG cutoff $\Lambda$ decreases, a divergent flow indicates an instability towards a symmetry-broken state in either the particle-particle or particle-hole channel. Further details of the formalism are provided in the SM.
@@
We begin by examining the electronic instability of the NiO2 bilayer at the interface. Fig. 2(a) presents the orbital-resolved band structure from the effective TB model for this bilayer, and Fig. 2(b) illustrates the corresponding Fermi surfaces. The onsite energy difference for the $d_{z^{2}}$ orbital between the top and bottom layers is approximately 0.23 eV according to Wannier fitting. Compared to the pressurized bulk case, there is an increase in the weight of both the $d_{x^{2}-y^{2}}$ bonding orbital on the $\alpha$ pocket and the $d_{z^{2}}$ antibonding orbital on the $\beta$ pocket. This occurs as the $d_{z^{2}}$ antibonding orbital moves away from the Fermi level, while the antibonding orbital shifts closer to it. In Fig. 2(c), we depict the representative FRG flow for a slightly hole-doped case with a reasonable interaction setting of $U=3$ eV and $J/U=0.1$. The SDW state is dominant from the outset and diverges at a relatively high cutoff, indicating a strong SDW instability, while superconductivity instability is weak. The inset shows the leading eigenvalues in the 2D Brillouin zone (BZ), with divergence occurring at $\bm{q}_{4}\approx(3\pi/4,3\pi/4)$, corresponding to Fermi surface nesting between the $\alpha$ and $\beta$ pockets. To simulate charge doping in realistic systems Zhou et al. (2025), we performed systematic calculations for both hole-doped ($\delta>0$) and electron-doped ($\delta<0$) cases for various Hund’s coupling. The resulting phase diagram, shown in Fig. 2(d), demonstrates a clear dominance of the SDW state with a wavevector on the $\Gamma$-M line across a wide doping regime. In contrast, superconductivity only appears within a narrow range of weak Hund’s coupling with small critical cutoffs. Notably, the leading pairing is $d$-wave for the electron-doped case after the Lifshitz transition (denoted by the black dashed line), where an electron pocket around the $\Gamma$ point arises from the $d_{z^{2}}$ antibonding orbital. Based on these findings, we deduce that the observed superconductivity is unlikely to originate from the interface, consistent with robust superconductivity observed in different samples with different interfaces in experiments Ko et al. (2025); Zhou et al. (2025); Liu et al. (2025).
@@
Figure 3: Band structure, Fermi surfaces and electronic instability for the NiO2 bilayer near the interface. Tight-binding model band structure (a) of the third NiO2 bilayer and corresponding Fermi surface (b) at filling $n=2.7$. (c) Typical FRG flows of leading eigenvalues in SC, SDW and CDW channels with $U=3$ eV, $J=0.1U,n=2.7$ of which leading instability is a $s_{\pm}$-superconductivity with inset figure showing the corresponding superconducting gap on Fermi surfaces. (d) Phase diagram of instability as a function of Hund’s rule coupling strength $J_{H}$ and hole doping level $\delta$ with a fixed $U=3$ eV. The symbols and color bar have the same meaning as in Fig.2. The black dashed line denotes the Lifshitz transition point $\delta_{c}=0.22$ with the appearance of hole pocket around the M point for $\delta>\delta_{c}$.

### v3

# Opposite-Mirror-Parity Scattering as the Origin of Superconductivity in Strained Bilayer Nickelates

@@ Abstract
We study the electronic structure and doping-dependent instabilities of strained La3Ni2O7 thin films using first-principles and functional renormalization group methods. We demonstrate that ordering tendencies are governed by Fermi surface scattering between electrons of opposite mirror parity. Under moderate hole doping, when the $d_{z^{2}}$ bonding band becomes incipient or crosses the Fermi level, robust $s_{\pm}$-wave superconductivity emerges from cooperative interlayer pairing reinforced by two competing spin-density-wave fluctuations. Compressive strain favors superconductivity in NiO2 bilayers slightly away from the interface, whereas tensile strain induces pair-breaking nesting that suppresses pairing. Our results establish a unified microscopic scenario for superconductivity in pressurized bulk and strained thin-film nickelates, providing new insights into high-Tc pairing in correlated quantum materials.
@@
Despite decades of extensive research efforts into nickelates, the nickel age of superconductors only began in 2019 with the discovery of superconductivity in ”infinite-layer” nickelates, specifically (Sr,Nd)NiO2 thin films grown on a substrate Li et al. (2019); Osada et al. (2020); Pan et al. (2022); Wang et al. (2022); Ding et al. (2023). Remarkably, four years later, nickelate superconductors entered the high-temperature regime akin to cuprates with the observation of superconductivity in a novel type of bilayer nickelate La3Ni2O7 (LNO) within the Ruddlesden-Popper phase under pressure, boasting an extraordinarily high transition temperature (Tc) of approximately 80 K Sun et al. (2023). Superconductivity emerges abruptly following a structural transition from the Amam phase to the Fmmm/I4mmm phase as pressure increases. In contrast to the $d^{9}$ electronic configuration of cuprates and ”infinite-layer” nickelates, the Ni2.5+ in LNO exhibits a $d^{7.5}$ configuration, with both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals contributing to the low-energy electronic structure due to the presence of apical oxygens between layers Luo et al. (2023); Zhang et al. (2023a); Lechermann et al. (2023a); Sakakibara et al. (2024a); Gu et al. (2025); Yang et al. (2024); Liu et al. (2024). Additionally, the trilayer nickelate La4Ni3O10 exhibits superconductivity under pressure with a lower Tc of 20-30 K Sakakibara et al. (2024b); Li et al. (2024a); Zhu et al. (2024); Zhang et al. (2025); Li et al. (2024b). Both bilayer and trilayer nickelates exhibit spin density wave (SDW) orders Sun et al. (2023); Zhang et al. (2024a); Liu et al. (2024); Chen et al. (2024); Ren et al. (2025), showing close relation with superconductivity Shi et al. (2025). The pairing mechanism and symmetry are under intensive theoretical study Yang et al. (2023a); Lu et al. (2024); Oh and Zhang (2023); Gu et al. (2025); Liu et al. (2023); Qu et al. (2024); Sakakibara et al. (2024a); Yang et al. (2023b); Qin and Yang (2023); Lu et al. (2023); Tian et al. (2024); Zhang et al. (2023b, 2024b); Jiang et al. (2024a); Lechermann et al. (2023b); Liao et al. (2023); Ryee et al. (2024); Luo et al. (2024); Fan et al. (2024); Jiang et al. (2024b, c); Zhan et al. (2025); Xia et al. (2025); Bejas et al. (2025); Duan et al. (2025). Theoretical calculations suggest that the pressure-driven Lifshitz transition, where the $d_{z^{2}}$ bonding band crosses the Fermi level ($E_{\text{F}}$) to form a hole pocket, is crucial for bulk superconductivity Sun et al. (2023). Spin fluctuations then promote an $s_{\pm}$-wave interlayer pairing Sakakibara et al. (2024a); Gu et al. (2025); Yang et al. (2023a); Liu et al. (2023). Other studies emphasize interlayer exchange, Hund’s coupling, and orbital hybridization in generating interlayer pairing in the strong-correlation limit Lu et al. (2024); Oh and Zhang (2023); Qu et al. (2024). However, further experimental investigation into the electronic structures and pairing gaps in the bulk nickelates is challenged due to the high-pressure conditions.
@@
Intriguingly, recent studies have shown that compressively strained LNO thin films grown on the SrLaAlO4 (SLAO) substrate exhibit superconductivity with a Tc exceeding 40 K at ambient pressure Ko et al. (2025); Zhou et al. (2025); Liu et al. (2025); Bhatt et al. (2025). Although this Tc is only about half the bulk value, these films provides a tunable and powerful platform for investigating the mechanisms underlying high-Tc superconductivity. First-principles calculations indicate that compressive strain eliminates the hole pocket around the M point, whereas tensile strain results in Fermi surfaces similar to those of the bulk material under pressure Geisler et al. (2025); Zhao and Botana (2025). This makes the absence of superconductivity in tensile-strained films Ko et al. (2025); Bhatt et al. (2025) a puzzling contradiction to the proposed bulk mechanism. Moreover, ARPES results on compressively strained LNO thin films are not yet consistent: one group reported heavy hole doping and a diffusive hole pocket around the M point Li et al. (2025); Yue et al. (2025); Shao et al. (2025), another identified that the $d_{z^{2}}$ bonding band resides slightly below $E_{\text{F}}$ Wang et al. (2025). These findings motivate a comprehensive study of key ingredients for superconductivity in bilayer nickelates and prompt three questions: (1) What mechanism drives superconductivity in strained films, and how does it related to the $d_{z^{2}}$ bonding band? (2) Why do tensile-strained LNO thin films, despite hosting Fermi surfaces similar to the bulk, fail to exhibit superconductivity? (3) Do bulk and thin film LNO share the same pairing mechanism?
To address these questions, we examine the electronic structure and doping-dependent correlated instabilities of strained LNO thin films by employing first-principles and functional renormalization group (FRG) calculations. We uncover an intimate relation between ordering tendencies and Fermiology, summarized in Fig. 1(a), where scattering between Fermi surfaces of opposite mirror parity (interlayer bonding/antibonding) plays an essential role in promoting correlated states. In compressively strained undoped films, nesting between the bonding $\alpha$ and antibonding $\beta$ pockets results an SDW with interlayer antiferromagnetic (AFM) configuration. With hole doping, the $d_{z^{2}}$ bonding band moves toward $E_{\text{F}}$; once it becomes incipient, additional finite-energy spin fluctuations arise from nesting between the antibonding $\beta$ pocket and the flat $d_{z^{2}}$ bonding band. These two SDW fluctuations, with different in-plane wave vectors, closely compete and suppress the SDW instability. However, they cooperatively promote interlayer pairing, yielding robust $s_{\pm}$-wave superconductivity, as shown in Fig. 1(b). When a bonding $\gamma$ pocket fully appears, $\beta$$\gamma$ nesting enhances interlayer pairing, whereas $\alpha$$\gamma$ and $\gamma$$\gamma$ nestings generate ferromagnetic interlayer coupling and suppress interlayer pairing. The latter effect is stronger, causing $T_{c}$ to decrease and resulting a Tc maximum near the Lifshitz transition. Our calculations suggest that superconductivity most likely originates from the NiO2 bilayer slightly away from the interface—where the $d_{z^{2}}$ bonding band resides near $E_{\text{F}}$ and substrate-induced hole doping can occur—rather than from the interfacial bilayer. Under tensile strain, the $\alpha$$\gamma$ nesting is significantly enhanced, weakening interlayer pairing and driving an SDW instability despite the presence of the $\gamma$ pocket. Finally, we discuss a unified scenario for superconductivity in both bulk and thin films and explore potential experimental implications.
Figure 1: (a) Schematic temperature-doping phase diagram for compressively strained LNO thin films. Insets show three representative Fermi surfaces together with nesting vectors: blue arrows indicate nesting between bonding (red) and antibonding (blue) pockets, while the orange arrow indicates nesting between two bonding pockets. The incipient $d_{z^{2}}$ band is represented by the shaded sector around the corner and $\delta_{c}$ marks the Lifshitz transition. (b) Schematics for the SDW with the wavevector of $\bm{q}_{4}$ and superconductivity with interlayer pairing.
@@
Electronic structures of LNO/SLAO interface. We start with the crystal and electronic structures of three-unit-cell (3UC) pristine LNO on the SLAO substrate. The tetragonal LNO is assumed and initial interface is constructed according to the structure from scanning transmission electron microscope measurements Zhou et al. (2025); Yue et al. (2025) and further relaxation is done for the 3UC LNO and the first two layers of substrate. The details of first-principles calculations are shown in the supplementary materials (SM). Fig. 2(a) shows the optimized crystal structure, where only three bilayers of the film are shown along with part of the substrate, and the complete structure is available in SM. The in-plane compressive strain generates an elongation along the c-axis for LNO, which mainly occurs on the LaO layers according to our calculations. To clearly demonstrate different structural changes at and near the interface, we select the first and third bilayers as representative examples and show their structural parameters. The out-of-plane distance between two Ni sites $d_{\text{Ni}-\text{Ni}}$ and outer apical Ni-O bond length $d_{\text{Ni}-\text{O}}$ within these bilayers are indicated in gray and red numbers, respectively. Across these bilayers, $d_{\text{Ni}-\text{Ni}}$ is approximately 4 Å, close to the bulk value, and this value is slightly longer in the 1st bilayer than the 3rd bilayer due to the effect of substrate. Notably, the outer apical Ni-O distance is increased to 2.3 Å  away from the interface due to the elongation of LaO layers and $d_{\text{Ni}-\text{O}}$ in the 1st bilayer is elongated to 2.6 Å. This elongation of the outer apical Ni-O bond results in a significant reduction in the onsite energy of the $d_{z^{2}}$ orbital in the 1st bilayer. Besides structural modifications, there is electrostatic doping from the substrate. Fig. 2(b) displays charge density differences and blue and red indicate regions of charge accumulation and depletion, respectively, suggesting a net electron transfer from the substrate to the 1st bilayer.
@@
Furthermore, the orbital-resolved band structures along high-symmetry paths for the 1st and 3rd bilayers are presented in Fig. 2(c) and (d), with a focus on the $e_{g}$ orbitals ($d_{x^{2}-y^{2}}$ and $d_{z^{2}}$). In the interfacial (1st) bilayer, the reduced onsite energy of the $d_{z^{2}}$ orbital induces a significant downward shift of its associated bands, as seen in Fig. 2(d). This shift causes the $d_{z^{2}}$ bonding band to move away from $E_{\text{F}}$, while the $d_{z^{2}}$ anti-bonding states approach and make contact with $E_{\text{F}}$ at the $\Gamma$ point. The electron doping is approximately 0.34 without considering substrate-induced Sr diffusion. Consequently, the 1st bilayer hosts a single $\beta$ hole pocket around the M point from a mixture of $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals, and an $\alpha$ electron pocket around the $\Gamma$ point from the $d_{x^{2}-y^{2}}$ orbital. The asymmetry between two Ni sites splits the $d_{x^{2}-y^{2}}$ bands along the $\Gamma-\text{M}$ path. Incorporating interaction in the LDA+U formalism will induce an additional electron pocket around $\Gamma$ from the $d_{z^{2}}$ antibonding state (see SM). In contrast, the bilayer away from the interface, such as the 2nd and 3rd bilayers, show no charge doping and have band structure similar to bulk LNO at ambient pressure, as shown in Fig. 2(c). The $d_{z^{2}}$ bonding band just drops below $E_{\text{F}}$  Zhao and Botana (2025), yielding an electron pocket around the $\Gamma$ point and a hole pocket around the M point. The substrate-induced Sr diffusion Zhou et al. (2025) will induce hole doping and make $d_{z^{2}}$ bonding band cross $E_{\text{F}}$ in these bilayers (see SM).
@@
Scenarios of superconductivity in compressed LNO thin films. Although superconductivity has been observed in strained LNO ultrathin films, its origin and which NiO2 bilayer hosts it remain elusive. Pinpointing the active bilayer is crucial for revealing the mechanism, as interfacial and interior bilayers have markedly different electronic structures. To investigate this issue, we construct effective tight-binding models by fitting hopping parameters to ab initio band structures and study electronic instability by varying carrier doping and interaction parameters, based on FRG calculations. As discussed earlier, the Ni $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals dominate the low-energy physics in strained LNO thin films, similar to bulk LNO at both ambient and high pressures Luo et al. (2023); Zhang et al. (2023a); Lechermann et al. (2023a); Sakakibara et al. (2024a); Gu et al. (2025); Yang et al. (2024); Liu et al. (2024). Therefore, a two-orbital tight-binding (TB) model can accurately describe the low-energy band structure Luo et al. (2023); Zhang et al. (2023a); Lechermann et al. (2023a); Sakakibara et al. (2024a); Gu et al. (2025), and the corresponding Hamiltonian is given by,
@@
Here $i,j=(m,l)$ label the in-plane lattice site ($m$) and top/bottom layer index ($l=t,b$), $\sigma$ labels spin, and $\alpha,\beta=x,z$ represent the Ni orbitals with $x$ denoting the $d_{x^{2}-y^{2}}$ and $z$ the $d_{z^{2}}$ orbital. The chemical potential $\mu$ varies with charge doping. Details of the Hamiltonian matrix are provided in the SM. Compared to the bulk case, the asymmetry between the two Ni sites in the 1st bilayer at the interface breaks the mirror symmetry connecting them, resulting in slightly different onsite energies and in-plane hopping for the two layers. The bilayer slightly away from the interface preserves this mirror symmetry and shares the same TB model with the bulk, though with modified effective parameters. For the electronic interactions, we adopt the general multi-orbital Hubbard interactions,
@@
where $U/U^{\prime}$ is the intra-/inter-orbital Hubbard repulsion, $J_{H}$ is the Hund ’s coupling, and $J_{P}$ is the pair-hopping interaction. The standard Kanamori relations $U=U^{\prime}+2J_{H}$ and $J_{H}=J_{P}$ are adopted in this work Castellani et al. (1978). To resolve electronic instabilities, we employ the FRG approach, which treats all particle-hole and particle-particle channels on equal footing, providing an unbiased depiction of correlated states from weak to intermediate coupling regimes Wang et al. (2009); Metzner et al. (2012); Platt et al. (2013). As the RG cutoff $\Lambda$ decreases, a divergent flow indicates an instability towards a symmetry-broken state in either particle-particle or particle-hole channel. Further details of the formalism are provided in the SM.
@@
We begin by examining the electronic instability of the interfacial NiO2 bilayer. Compared to the pressurized bulk case, there is an increase in the weight of both the $d_{x^{2}-y^{2}}$ bonding orbital on the $\alpha$ pocket and the $d_{z^{2}}$ antibonding orbital on the $\beta$ pocket. Systematic calculations for both hole-doped ($\delta>0$) and electron-doped ($\delta<0$) cases across various Hund’s couplings show that the phase diagram is predominantly governed by a spin-density-wave (SDW) state associated with Fermi surface nesting between the $\alpha$ and $\beta$ pockets. In contrast, superconductivity only appears within a narrow range of weak Hund’s coupling with low critical cutoffs. These findings suggest that superconductivity is unlikely to originate from the interfacial NiO2 bilayer, consistent with robust superconductivity observed in samples with different interfaces Ko et al. (2025); Zhou et al. (2025); Liu et al. (2025).
@@
Figure 3: (a) TB-model band structure of the third NiO2 bilayer and corresponding Fermi surface (b) at the filling $n=2.7$ marked by the green dashed line in (a). (c) Typical FRG flows of leading eigenvalues in SC, SDW and CDW channels with $U=3$ eV, $J=0.1U,\delta=0.3$. Superconductivity is leading with two SDW competing and sub leading and the corresponding superconducting gap is shown in the inset. (d) Phase diagram of instabilities as a function of Hund’s coupling $J_{H}$ and hole doping $\delta$ at fixed $U=3$ eV. The color bar indicates the critical scale $\Lambda_{c}$ and the black dashed line denotes the Lifshitz transition point $\delta_{c}=0.22$.